IODP Proceedings    Volume contents     Search


Site M00271

Expedition 313 Scientists2


Mobilization of the L/B Kayd

Mobilization of the L/B Kayd began with shipping the European Consortium for Ocean Research Drilling (ECORD) Science Operator (ESO) laboratory, database, and office containers from ESO partner institutes in Europe. Once all the equipment and containers had cleared customs, they were delivered to the U.S. Coast Guard Station in Atlantic City, New Jersey (USA).

On 24 April 2009, the L/B Kayd arrived ahead of schedule at the Coast Guard Station in Atlantic City. Mobilization was started immediately by the drilling contractor, Drilling, Observation and Sampling of the Earth's Continental Crust (DOSECC), who loaded the drilling and ESO containers onto the platform. ESO staff from the British Geological Survey (BGS) arrived at the L/B Kayd on 28 April to prepare the containers for operations, including installing a shipboard computer network. Because of the nature of transiting in a lift boat, with water spilling over the deck, power could not be fully installed in the containers until the first hole was reached.

Mobilization continued for 6 days. The wireline logging equipment was transferred to the L/B Kayd out of Miller's Launch, Staten Island, New York (USA), on 1 May, with the first supply boat and crew change.

Transit to Hole M0027A

At 1030 h Universal Time Coordinated (UTC) on 30 April 2009, the L/B Kayd set sail from Atlantic City and headed for the first coring site (Integrated Ocean Drilling Program [IODP] Hole M0027A). A brief stop was made on the way to Hole M0027A to modify the flooring of the cantilevered drilling platform, which was being forced up by waves breaking on the L/B Kayd's bow. At 2345 h, the L/B Kayd arrived on site and prepared to position above Hole M0027A.

Hole M0027A

At 0003 h on 1 May 2009, the L/B Kayd was positioned above Hole M0027A and the legs were lowered to tag the seabed. Once the seabed was tagged, the preload procedure began by gradually increasing the load on the seabed. The preload procedure was interrupted to reposition the L/B Kayd so that the communication satellites were not eclipsed by the legs. The preload and settlement procedure continued until 1053 h, when the L/B Kayd was jacked-up to ~30 ft above the water. Normal access to the working deck was granted, and all teams prepared to begin coring. At 1215 h, the supply vessel Rana Miller arrived at the platform and delivered equipment and personnel, which included the wireline logging tools and the remaining scientific staff. At 1640 h, the drill rig was started, the mast was raised, and the conductor pipe was run to the seabed, tagging it at 2134 h. The core barrel was lowered to just above the seabed, and the first hydraulic piston corer (HPC) core was fired at 2350 h (Table T1).

The first core arrived on deck at 0010 h on 2 May. Coring continued using the HPC, and the hole was advanced by recovery. Sixteen HPC attempts resulted in two nonfires and a penetration of 28 m in just under 20 h. Three core runs showed signs of caving, with the amount of caving fill between 0.33 and 0.72 m. Once the base of the sand layer had been established, coring was stopped. At 2035 h, the drill string was tripped, and the conductor pipe was extended into the seabed to act as casing.

At 0130 h on 3 May, the conductor pipe was set at 17.6 m drilling depth below seafloor (DSF). The hole was conditioned, and the PQ drill string was run to the base of the hole, where HPC coring commenced. After two HPC attempts, the coring tool was switched to the extended nose corer (EXN) because of a hard lithology preventing HPC recovery. However, recovery was still poor. The coring method was switched back to the HPC, after which 6 m of cavings was encountered. In an attempt to stabilize the hole, the mud was changed and the hole was reamed with a full-face bit in a standard rotary corer (ALN) barrel. Hole stability problems continued, and the full-face bit was found to have no cutting surface left when retrieved at 1830 h. The string was tripped back to the surface to inspect the outer bit; however, no damage was observed. The string was rerun to the base of the hole, and coring continued with the HPC, which, when retrieved, contained a few pebbles and had a dented nose. The ALN was run to clear the hole shortly before midnight.

During the rerun of the ALN corer, the bit blocked again. A second ALN corer was run at 0100 h on 4 May, which on the second attempt advanced the full 3 m but recovered no core. At 0353 h, HPC coring advanced the hole but ended when the drill string became stuck in the hole. After the drill string was freed, it was recovered to the deck and checked for damage. The string was lowered back into the hole to within 9 m of the base, where it was stopped by infill material. It was also suspected that pieces of the previously broken full-face bit had been encountered, and the hole was reamed out with the ALN corer to save the PQ string bit. Coring recommenced using the EXN tool. At 1620 h, it was decided that the depth of casing in the hole should be increased. After pulling the PQ string, the casing could not be moved in the hole, so the hole was reentered using the PQ string and the hole was advanced using the EXN tool. The EXN was used in favor of the HPC, as the HPC seemed to increase hole collapse.

By 0400 h on 5 May, Hole M0027A was good, clean, and free-running with circulation at the base. EXN coring continued throughout the day with mixed recovery, although infill was encountered with most new core runs. The core barrel handling procedure was modified so that two EXN core barrels were always operational, which reduced the time the hole was left vulnerable to cave-ins. This, coupled with using the EXN, dramatically improved the coring. From 2000 h, an increasingly hard lithology caused core runs to become blocked, and a liner was crushed at the shoe.

At 2355 h on 5 May, operations were halted because of a thunderstorm.

Coring restarted at 0255 h on 6 May using the ALN, but problems with crushed liners blocking the bit persisted throughout the first half of the day. At 1500 h, the mud mix was altered slightly, and the ALN corer was switched to the HPC tool, which improved recovery slightly. At 1830 h, a meeting was convened to discuss how to best advance the hole, as progress to date had been slow. All agreed to HPC spot core in maximum 60 ft increments (6 × 10 ft pipe lengths, ~18 m), or sooner if a variation in drilling parameters was encountered that suggested a change in lithology. This was done to ~180 m DSF, where continuous coring resumed. The first open-hole interval commenced at 2140 h and continued past midnight.

Rapid progress was made at the beginning of 7 May using the strategy of HPC spot coring every 60 ft or less. Seven HPC attempts over 64 m were made in 20 h, with a slight setback occurring when the HPC tool became stuck in the bottom-hole assembly (BHA) and the wireline wire broke during an attempt to pull the tool free. The result of this was that the PQ string had to be pulled out of the hole, which was a wet pipe trip because of the HPC sealing the BHA. The wireline was not fished, as there was no way to circulate the mud during fishing, which would have increased the risk of hole collapse.

The wireline wire, which had broken ~80 ft above the BHA, was replaced early in the morning of 8 May, as was a worn bit on the BHA. The drill string was tripped back in the hole with the noncoring bit inserted. The string reached the base of the hole 20 m higher than expected because of infill, which was subsequently drilled out. The noncoring barrel got stuck 3 m above the base of the hole when it was unlatched from the BHA. This necessitated another drill string trip. After 14 h, HPC coring resumed at 1325 h. The liner collapsed in several runs because of very stiff clay, which limited core recovery and caused several core barrels to become stuck in the BHA.

Problems with the HPC tool getting stuck in the BHA continued for the first 3 h of 9 May. After partly tripping the drill string, the HPC became free and was returned to the deck. From 0330 h, EXN and HPC coring were alternately conducted, with the EXN tool eventually chosen in favor of the HPC tool. Seven runs with the EXN tool collected very good core before chattering on the drill string, and a lack of penetration indicated a change back to sand.

To tackle this lithology, the ALN corer was prepared, and early on 10 May coring resumed with the ALN in what appeared to be a gravel layer. There was no recovery in this layer. A mud backflow occurred, and large bubbles were observed rising up the drill string. No H2S or abnormal smell was recorded by the DOSECC or ESO instruments, and it was assumed that the gravel layer was hosting a freshwater flow. The next core did have an abnormal smell, and tests using the ESO gas analyzer measured 7.7 ppm H2S and no flammables. A later core tested positive (1%) for flammables near a thin very dark layer. Excellent coring conditions prevailed into 11 May through alternating hard clays and hard fine sands.

Progress slowed slightly when the lithology became more sandy until the hole collapsed and the drill string became stuck at 1150 h on 12 May. Upon freeing the string, 9 m of infill blocked the hole and an increase in rotational torque was observed. To reduce this pressure on the string, thought to be from the clay formation above, the string was pulled back a total of 26 m to ream the hole. Progress continued through very loose, clean sand for the rest of the day, with surprisingly excellent core recovery. On 13 May, very good progress continued using the ALN corer, reaching 451.06 m DSF at midnight.

By 0430 h on 14 May, it became apparent that the BHA stabilizer rings were worn and needed replacement. The string was tripped in order to replace both rings. The trip was completed by 0800 h, the rings replaced, and the pipe run back into the hole. The bottom of the hole was tagged at 335 m DSF because of a bridge that had formed, at which depth open holing commenced to clear it. A zone of high-pressure water was encountered, which continued until 345 m DSF. Below this, the backpressure decreased to normal, and the string broke through the blockage. Infill was encountered after another 8–10 m air gap beneath the bridge. Approximately 100 m of fill was drilled out to reach the bottom of the hole. Coring recommenced using the ALN corer at 1910 h.

Progress slowed on 15 May, although core recovery was generally good with near 100% for many of the runs. However, some sections of some cores were undersized. The ALN corer was replaced by the EXN corer for one run to see if it gave better results. The lithology proved to be unsuitable for EXN coring with little penetration, high backpressure, and torque, so the ALN corer was used for the rest of the day. In the evening, the core barrel got stuck in the BHA at 509 m DSF. It was eventually freed, but after this event, core recovery was poor and there were ongoing problems with core barrel latching and retrieval. By the end of the day, the depth of the hole was 515 m DSF.

Steady progress was made on 16 May with the ALN corer. Core slippage occurred on some core runs, but the core was usually retrieved on the next run. Good progress was made on 17 May, with most cores having 100% recovery. Repairs to the wireline and bit refurbishment caused some delays in the afternoon and early evening.

On 18 May, 11 core runs were made with >30 m penetration using the ALN corer. Recovery was more variable than in recent days (85%). In some cases this was due to core slippage, although slipped material was often recovered in the next run. The rate of penetration slowed because of a combination of the formation becoming harder and the round trip of the core barrel taking longer with increasing depth in the hole. There were reports of a "petroleum smell" from the drillers in the early morning. However, there was no reading on the gas monitors. A slight increase in torque was noticed in the evening.

Shortly before midnight on 18 May, the core barrel was recovered to deck with no core. It was suspected that the core barrel had not latched into the BHA. Another barrel was deployed and became jammed in the upper section of the drill string. This was discovered when the overshot latched onto the barrel at approximately the waterline. After removing 11 double stands of pipe, the overshot with the latch head was unable to pass through, although the overshot alone could. The core barrel was deployed with the latching indicator ball installed and successfully latched into the BHA. After adding pipe (0410 h, 19 May) to return the string to the base of the hole, increased torque, backlash during rotation, and backpressure on the mud gauge was noted. When the string was broken to add further lengths, frothing and an overflow of mud occurred. On lowering the overshot, slack on the wireline indicated it was being temporarily stopped in the drill string. The overshot latched on as normal, but on recovery, the assembly became jammed at ~25 m below deck level. After two attempts to release the core barrel, it was recovered to deck. Scratches and spiral polishing were noted on the outside of the barrel.

At 0600 h on 19 May, the coring operation in Hole M0027A was stopped (base of last core; 631 m DSF) and preparations made to start the logging program. Because of the instability in the upper part of the hole, the decision was made to log the open hole in three sections. Through-pipe total gamma ray (TGR) was acquired along almost the full length of the hole. Open-hole resistivity, magnetic susceptibility, sonic velocity, and spectral gamma ray were acquired for the bottom section from 421 to 631 m DSF. Acoustic images were acquired in the top half of this lowermost interval. The pipe was then pulled to 195 m DSF, beginning at 1455 h on 20 May, and resistivity, magnetic susceptibility, spectral gamma ray, sonic velocity, and acoustic image logs were acquired between 195 and 342 m DSF. A bridge that had formed at 342 m DSF prevented logging of the section between 342 and 421 m DSF.

At dawn on 21 May, preparations were made to begin a vertical seismic profile (VSP), beginning with a marine mammal watch at 0513 h. Air gun firing began at 0548 h, and the VSP tool was inserted into the pipe at 0600 h. VSP work continued throughout the day until ~1900 h. Open-hole VSP data were acquired from 323 to 195 m DSF with through-pipe VSP data from 195 m DSF to seafloor. Marine mammal observations ceased at 1930 h.

Preparations were made to pull the pipe up to the top of the next open-hole logging section. However, a fault developed with the drilling rig motor before pipe pulling commenced. Repairs required a spare part that was not on the platform. Efforts to make a temporary repair on board to keep the system operational for the remainder of the logging program were attempted but were unsuccessful. Late in the evening, operations ceased while waiting for the part (and additional spare) to be delivered by supply boat on 22 May.

The rig was operational by 0830 h on 22 May. Two double stands were tripped with the aim of opening the hole to 97 m DSF. However, the pipe became stuck, and despite applying five times the usual torque, the pipe would not rotate in the hole and could not be pulled further. The decision was made to abandon the hole and move to Hole M0028A to begin coring operations. The pipe was cut just above the core barrel and tripped. The casing was retrieved, and the deck, drilling rig, and containers were secured for the rig move.

1Expedition 313 Scientists, 2010. Site M0027. In Mountain, G., Proust, J.-N., McInroy, D., Cotterill, C., and the Expedition 313 Scientists, Proc. IODP, 313: Tokyo (Integrated Ocean Drilling Program Management International, Inc.). doi:10.2204/iodp.proc.313.103.2010

2Expedition 313 Scientists' addresses.

Publication: 4 December 2010
MS 313-103